OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (zpupster):

you are in the 4th quadrant

OpenStudy (zpupster):

 tan(θ) = r sin(θ) / (r cos(θ)) = -2 / 2 = -1

OpenStudy (zpupster):

tan^-1 = 315 315 - 180 is the other angle

OpenStudy (zpupster):

2 = r cos(45) solve for this to get r (r,theta)

OpenStudy (anonymous):

thank you